OpenStudy (anonymous):

Solve 2x-8/4 = 5. Show your work.

OpenStudy (anonymous):

is it (2x-8)/4? or2x-(8/4)

OpenStudy (anonymous):

2x - 8 / 4 appears like a fractin.

OpenStudy (anonymous):

If as problem follows, then 8/4 is first, which equals 2. Then 2x-2=5 +2 on both sides to isolate x 2x=7 divide by 2 on both sides x=7/2 or 3 and 1/2

OpenStudy (anonymous):



OpenStudy (anonymous):

I know the answer is 14 I just don't know how to get to the answer.

OpenStudy (anonymous):

ohhhh so (2x-8)/4=5 multiply by 4 on both sides to isolate x 5 times 4 = 20 2x-8=20 plus 8 on both sides 20+8=28 2x=28 divide by 2 on both sides 28/2=14 x=14

OpenStudy (anonymous):

Thanks (:

OpenStudy (anonymous):

np